Find the mean of the given probability distribution. 15) 15) The number of golf balls ordered by customers of a pro shop has the following probability distribution. 3 0.14 6 0.25 90.36 12 0.15 15 0.10
Answer #1

Solution :

From given probability distribution,

Mean = \mu = \sum X * P(X) =

= 3 * 0.14 + 6 * 0.25 + 9 * 0.36 + 12 * 0.15 + 15 * 0.10

= 8.46

Mean = \mu = 8.46